# Feedback Linearisation with State Constraints

### v1

@@ 3.1 Ensuring an Integrable Input
Given a system $\Sigma_{A}=(f_{A},g_{A},h_{A},x_{A}(t_{0}))$ (Eq. (14)) we define the following system integral controller map
@@ 4.1 FBL with Pole Placement
FBL converts nonlinear dynamics into linear dynamics by employing appropriate feedback control. Considering a system represented by Eq. (4) and a reference signal, $y_{r}\in C^{\infty}([t_{0},\infty),\mathbb{R}^{m})$, we next recall the standard FBL approach that derives an expression that relates the tracking error $E(t):=y(t)-y_{r}(t)$ to the system input by taking successive time derivatives of tracking error. Having obtained this expression, we can design a feedback law to cancel out any nonlinearity as well as terms not involving the tracking error. The remaining system will be linear allowing us to utilise linear controllers for stabilisation to ensure that tracking errors asymptotically approach zero. In this paper, we choose the pole-placement strategy as it is frequently used for linear system stabilisation, however, it does not mean that the pole-placement is the only standard strategy.
@@
$\begin{bmatrix}E_{1}^{(\sigma_{i})}(t)\\ \vdots\\ E_{m}^{(\sigma_{i})}(t)\end{bmatrix}=\Gamma_{f}(t,x(t))+\Gamma_{g}(t,x(t))u(t)-\Gamma_{r}(t).$ (28)
@@
is inevitable for all $(t,x)\in[t_{0},\infty)\times\mathbb{R}^{n}$.
@@
we simply expand $(\lambda-\bar{\lambda}_{i,1})\dots(\lambda-\bar{\lambda}_{i,\sigma_{i}})$ and equate with the characteristic equation, $\det(\lambda I-(A_{i}-BK_{i}))$. By equating the coefficients of the same powers of $\lambda$, we solve for the $K_{i,j}$’s, selecting the appropriate gains for this pole placement.
If the real parts of the poles of a linear system are selected as negative, then it is well known that the system is exponentially stable. In the context of ODE (30), this means that the tracking error, as well as the tracking error time derivatives, exponentially tend to zero. Poles can be adjusted so that criteria such as settling time, percentage overshoot, etc, are within a specified range.
@@ 4.2 Challenges: Augmentation with FBL
Even when $\Omega_{g}$ or $\Gamma_{g_{I}}$ remains theoretically invertible, they may be difficult to invert under finite precision arithmetic computation if their values approach zero, leading to numerical instability. However, since both augmentation and FBL only require invertibility of $\Omega_{g}$ and $\Gamma_{g_{I}}$ along the closed-loop trajectory, it may still be feasible to apply these methods provided that the trajectory avoids regions where these matrices are non-invertibile. If the system approaches a non-invertibility region, increasing $\beta$ can help mitigate ill-conditioning of $\Gamma_{g_{I}}$, as Lemma 2 shows that $\Gamma_{g_{I}}$ depends linearly on $\beta$. In cases where the system trajectory passes directly through singular regions, we introduce in Section 5 a switching strategy that adapts to the system’s local relative degree.
@@ 5 A Practical Approach for FBL with State Constraints
The key idea is to synthesise different closed-form controllers for regions with varying “local" relative degrees and switch between them as the trajectory transitions through these regions. Following [32, 33], we restrict our focus to SISO systems ($m=1$), where $\Omega_{g}(t,x):=L_{g}L_{f}^{\rho_{1}-1}\phi_{1}(t,x)$ and $\Gamma_{g}(t,x)=L_{g}L_{f}^{\sigma-1}h(t,x)$ are both scalar. For brevity, we refer to [32, 33] for conditions ensuring asymptotic tracking for switching FBL controllers. This work focuses on implementing the switching strategy in the context of our state constrained FBL framework, where we apply a localised relative degree (Definition 4) to augment the system, capturing constraints, and perform FBL controller synthesis.
@@ 5.1 Constraint Capture Using $\varepsilon$-Numerical Relative Degree
$\frac{d^{\hat{\rho}_{k}}}{dt^{\hat{\rho}_{k}}}\left(\phi_{k}(t,x(t))+\frac{1}{2}z^{2}_{k}(t)\right)\approx\frac{1}{2}\sum_{j=1}^{\hat{\rho}_{k}}{\hat{\rho}_{k}\choose j}z_{k}^{(\hat{\rho}_{k}-j)}(t)z_{k}^{(j)}(t)$
$\qquad+z_{k}(t)z_{k}^{(\hat{\rho}_{k})}(t)+L_{g}L_{f}^{\hat{\rho}_{k}-1}\phi_{k}(t,x)u(t)+L_{f}^{\hat{\rho}_{k}}\phi_{k}(t,x).$ (33)
@@
$\omega_{k,(t_{0},x_{0})}(t,x):=\left[\lvert{L_{g}L_{f}^{0}\phi_{k}(t,x)}\rvert,...,\lvert{L_{g}L_{f}^{\hat{\rho}_{k}(t_{0},x_{0})-1}\phi_{k}(t,x)}\rvert\right$] (38)
@@
Input: System $\Sigma$ (3), Constraints $\phi$ (25), Threshold $\varepsilon$,
@@ 5.2 FBL with $\varepsilon$-Numerical Relative Degree
We use $\omega_{k,(t_{0},x_{0})}(t,x)$ (Eq. (38)) and $\gamma_{(t_{0},x_{0})}$ (Eq. (41)) to form switching sets to detect the change of $\varepsilon$-NRDs, with details presented in Section 5.3 with Eqs. (42)-(43). If the $\varepsilon$-NRD changes, then we must update Eq. (39) and synthesise a new FBL controller based on the current temporal and state space coordinates. We have summarised how to numerically synthesise such an FBL for a given state $(t_{0},x_{0})$ and threshold $\varepsilon$ for a general SISO system $\Sigma$ in Algorithm 2.
@@ 5.3 Controller Implementation
$C_{1}$ $:=\bigcap_{k=1}^{r}\Bigl{\{}(t,x,\xi):[\omega_{k,(t_{0},x_{0})}(t,[x,\xi])]_{i}<\epsilon_{i}$ (42)
 $\qquad\qquad\text{ for }1\leq i\leq\hat{\rho}_{k}(t_{0},x_{0})-1,\;$
$\text{and}\;[\omega_{k,(t_{0},x_{0})}(t,[x,\xi])]_{\hat{\rho}_{k}(t_{0},x_{0})}\geq\epsilon_{\hat{\rho}_{k}(t_{0},x_{0})}\Bigr{\}},$
$C_{2}$ $:=\Bigl{\{}(t,x,\xi):\gamma_{i}(t,[x,\xi])<\epsilon_{i}\text{ for }1\leq i\leq\hat{\sigma}(t_{0},x_{0})-1,\;$
$\qquad\qquad\text{and}\;\gamma_{\hat{\sigma}(t_{0},x_{0})}(t,[x,\xi])\geq\epsilon_{\hat{\sigma}(t_{0},x_{0})}\Bigr{\}}.$
@@ 6 Numerical Experiments
$\pi(t,[x,\xi])=\frac{-1}{s_{\beta}^{\prime}(\xi(t))\sqrt{16(t-0.5)^{2}-2x_{1}(t)-1}}$ (53)
$\qquad\times\bigg{(}3\sqrt{16(t-0.5)^{2}-2x_{1}(t)-1}s_{\beta}^{\prime}(\xi(t))$
$\qquad\qquad-K\bigg{[}x_{1}(t),x_{2}(t),16-\frac{(16t-x_{2}(t)-8)^{2}}{16(t-0.5)^{2}-2x_{1}(t)-1}$
$\qquad\qquad-s_{\beta}(\xi(t))\sqrt{16(t-0.5)^{2}-2x_{1}(t)-1}\bigg{]}^{\top}\bigg{)},$
@@
$u(t,[x,\xi])=x_{2}+16-\frac{(16t-x_{2}-8)^{2}}{16(t-0.5)^{2}-2x_{1}-1}$
$\hskip 71.13188pt-s_{\beta}(\xi)\sqrt{16(t-0.5)^{2}-2x_{1}-1},$ (54)
@@
To illustrate the influence of pole placement on controller performance, we performed numerical experiments using Algorithm 1 and Algorithm 2 to synthesise closed-loop controllers for different pole locations. The resulting simulations of system trajectories are shown in Fig. 2a. Each simulation applied the derived controller to the system while monitoring the switching sets from Eq. (42), which simplifies to Eq. (55) in this example. Whenever a change in the $\varepsilon$-NRD was detected, the controller was re-synthesised about the current state trajectory using the same procedure. The parameters were fixed as $\beta=100$ and $\varepsilon=[0.01,\dots,0.01]^{\top}$, with poles assigned either to 2.9-2.9 or 12-12.
@@
$\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\\ \dot{x}_{3}\end{bmatrix}$ $=\begin{bmatrix}10(x_{1}-x_{2})\\ 28x_{1}-x_{2}-x_{1}x_{3}\\ x_{1}x_{2}-\frac{8}{3}x_{3}\end{bmatrix}+\begin{bmatrix}0\\ 1\\ 0\end{bmatrix}u,\text{ }x(0)=\begin{bmatrix}0.1\\ 1\\ 15\end{bmatrix}$
$y(t)$ $=x_{2}(t)\text{ with }y_{r}(t)=0.6(\sin(1.5t)-\cos(1.3t))$
$\phi(t,x)$ $=[\phi_{1}(t,x),\phi_{2}(t,x)]^{\top}=[-x_{2}-1.2,x_{2}-1.2]^{\top}\leq 0.$
@@
Numerical experiments were conducted using an integral parameter of $\beta=20$, $\varepsilon$-NRD threshold parameter of $\varepsilon=[0.01,\dots,0.01]$ and positioning the poles at 0.3-0.3. By initialising the virtual input to have zero initial value according to Eq. (24), the initial condition of integral states was found to be $\xi(0)=[\xi_{1}(0),\xi_{2}(0)]^{\top}=[0.0143,-0.0388]^{\top}$. We execute Algorithm 1 and Algorithm 2 to synthesise a controller and simulate the closed-loop system. During simulation, we monitor the switching condition in Eq. (43). When triggered, we re-execute two algorithms to synthesise a new controller, switch to the updated system, and continue the simulation, again tracking the switching condition.

### v2

@@ 3.1 Ensuring an Integrable Input
Given a system $\Sigma_{A}=(f_{A},g_{A},h_{A},x_{A}(t_{0}))$ (Eq. (14)) and a integration bound $\beta$, we define the following system integral controller map
@@ 4 Feedback Linearisation
incorporating the augmented structure from the previous section, given in Eq. (14), to handle state constraints.
@@ 4.1 FBL with Pole Placement
Considering a system represented by Eq. (4) and a reference signal, $y_{r}\in C^{\infty}([t_{0},\infty),\mathbb{R}^{m})$, we next recall the standard FBL approach that derives an expression that relates the tracking error $E(t):=y(t)-y_{r}(t)$ to the system input by taking successive time derivatives of tracking error. Having obtained this expression, we can design a feedback law to cancel out any nonlinearity as well as terms not involving the tracking error. The remaining system will be linear allowing us to utilise linear controllers for stabilisation to ensure that tracking errors asymptotically approach zero. In this paper, we choose the pole-placement strategy as it is frequently used for linear system stabilisation, however, it does not mean that the pole-placement is the only standard strategy.
@@
$\begin{bmatrix}E_{1}^{(\sigma_{1})}(t)\\ \vdots\\ E_{m}^{(\sigma_{m})}(t)\end{bmatrix}=\Gamma_{f}(t,x(t))+\Gamma_{g}(t,x(t))u(t)-\Gamma_{r}(t).$ (28)
@@
is invertible for all $(t,x)\in[t_{0},\infty)\times\mathbb{R}^{n}$.
@@
we simply expand $(\lambda-\bar{\lambda}_{i,1})\dots(\lambda-\bar{\lambda}_{i,\sigma_{i}})$ and equate with the characteristic equation, $\det(\lambda I-(A_{i}-BK_{i}))$. By equating the coefficients of the same powers of $\lambda$, we solve for the $K_{i,j}$’s, selecting the appropriate gains for this pole placement. We will later discuss the selection criteria for poles in Section 5.4.
@@ 4.2 Challenges: Augmentation with FBL
Even when $\Omega_{g}$ or $\Gamma_{g_{I}}$ remains theoretically invertible, they may be difficult to invert under finite precision arithmetic computation if their values approach zero, leading to numerical instability. However, since both augmentation and FBL only require invertibility of $\Omega_{g}$ and $\Gamma_{g_{I}}$ along the closed-loop trajectory, it may still be feasible to apply these methods provided that the trajectory avoids regions where these matrices are non-invertible. If the system approaches a non-invertibility region, increasing $\beta$ can help mitigate ill-conditioning of $\Gamma_{g_{I}}$, as Lemma 2 shows that $\Gamma_{g_{I}}$ depends linearly on $\beta$. In cases where the system trajectory passes directly through singular regions, we introduce in Section 5 a switching strategy that adapts to the system’s local relative degree.
@@ 5 A Practical Approach for FBL with State Constraints
The key idea is to synthesise different closed-form controllers for regions with varying “local" relative degrees and switch between them as the trajectory transitions through these regions. Following [30, 31], we restrict our focus to SISO systems ($m=1$), where $\Omega_{g}(t,x):=L_{g}L_{f}^{\rho_{1}-1}\phi_{1}(t,x)$ and $\Gamma_{g}(t,x)=L_{g}L_{f}^{\sigma-1}h(t,x)$ are both scalar. For brevity, we refer to [30, 31] for conditions ensuring switching FBL controllers result in an asymptotic approximate tracking with a bounded tracking error based on the magnitude of the switching threshold. This work focuses on implementing the switching strategy in the context of our state constrained FBL framework, where we apply a localised relative degree (Definition 4) to augment the system, capturing constraints, and perform FBL controller synthesis. Note that we only consider SISO systems when the boundary singularity issue arises. If the singularity issue never occurs, our work prior to this section remains valid for MIMO systems, and a sole FBL controller is perpetually employed to guarantee an asymptotic tracking under the pole placement strategy.
@@ 5.1 Constraint Capture Using $\varepsilon$-Numerical Relative Degree
$\frac{d^{\hat{\rho}_{k}}}{dt^{\hat{\rho}_{k}}}\left(\phi_{k}(t,x(t))+\frac{1}{2}z^{2}_{k}(t)\right)\approx$ $\frac{1}{2}\sum_{j=1}^{\hat{\rho}_{k}}{\hat{\rho}_{k}\choose j}z_{k}^{(\hat{\rho}_{k}-j)}(t)z_{k}^{(j)}(t)+z_{k}(t)z_{k}^{(\hat{\rho}_{k})}(t)$
$+L_{g}L_{f}^{\hat{\rho}_{k}-1}\phi_{k}(t,x)u(t)+L_{f}^{\hat{\rho}_{k}}\phi_{k}(t,x).$ (33)
@@
$\omega_{k,(t_{0},x_{0})}(t,x):=\left[\lvert{L_{g}L_{f}^{0}\phi_{k}(t,x)}\rvert,\dots,\lvert{L_{g}L_{f}^{\hat{\rho}_{k}(t_{0},x_{0})-1}\phi_{k}(t,x)}\rvert\right$] (38)
@@
Input: System $\Sigma$ (3), Constraints $\phi$ (25), Threshold $\varepsilon$, Integration Bound $\beta$
@@ 5.2 FBL with $\varepsilon$-Numerical Relative Degree
We use $\omega_{k,(t_{0},x_{0})}(t,x)$ (Eq. (38)) and $\gamma_{(t_{0},x_{0})}$ (Eq. (41)) to form switching sets to detect the change of $\varepsilon$-NRDs, with details presented in Section 5.3 with Eqs. (5.3)-(43). If the $\varepsilon$-NRD changes, then we must update Eq. (39) and synthesise a new FBL controller based on the current temporal and state space coordinates. We have summarised how to numerically synthesise such an FBL for a given state $(t_{0},x_{0})$ and threshold $\varepsilon$ for a general SISO system $\Sigma$ in Algorithm 2.
@@ 5.3 Controller Implementation
$C_{1}$ $:=\bigcap_{k=1}^{r}\Bigl\{(t,x,\xi):[\omega_{k,(t_{0},x_{0})}(t,[x,\xi])]_{i}<\epsilon_{i}\text{ for }1\leq i\leq\hat{\rho}_{k}(t_{0},x_{0})-1,\;$
$\qquad\qquad\qquad\qquad\text{and}\;[\omega_{k,(t_{0},x_{0})}(t,[x,\xi])]_{\hat{\rho}_{k}(t_{0},x_{0})}\geq\epsilon_{\hat{\rho}_{k}(t_{0},x_{0})}\Bigr\},$ (42)
$C_{2}$ $:=\Bigl\{(t,x,\xi):[\gamma_{(t_{0},x_{0})}(t,[x,\xi])]_{i}<\epsilon_{i}\text{ for }1\leq i\leq\hat{\sigma}(t_{0},x_{0})-1,\;$
$\qquad\qquad\qquad\qquad\text{and}\;[\gamma_{(t_{0},x_{0}))}(t,[x,\xi])]_{\hat{\sigma}(t_{0},x_{0})}\geq\epsilon_{\hat{\sigma}(t_{0},x_{0})}\Bigr\}.$
@@
5.4 Selection of Controller Synthesis Parameters
Parameters threshold $\varepsilon$, the integration bound $\beta$, and the system poles $\bar{\lambda}$, are required for Algorithms 1 and 2. In practice, a grid search may be used to select the controller parameters that yield the desired performance. We next provide some guidelines on why you should not select these parameters to be too large or too small.
  1. 1.
    The switching region (the compliment of $C_{1}\cap C_{2}$ from Eq. (5.3)) is a function of the threshold $\varepsilon$. Increasing $\varepsilon$ shrinks $C_{1}\cap C_{2}$, which increases the size of the switching region. By the same logic, decreasing $\varepsilon$ reduces the size of the switching region. Selecting a small $\varepsilon$ threshold gives a small switching region, requiring a small numerical ODE step size to avoid inadvertently jumping too far in the direction of the vector field, causing constraint violation. Small numerical step size implies that the implementation of the controller on a physical system requires a high sensor refresh rate. On the other hand, selecting a $\varepsilon$-threshold that is too large results in significant approximation errors when evaluating $\varepsilon$-NRDs (Definition 4), and may eventually degrade the performance of the controller (see [30] that provides a bound on asymptotic tracking error depending on the magnitude of $\varepsilon$ threshold).
  2. 2.
    Lemma 2 demonstrates that the integration bound $\beta$ depend on $\Gamma_{g_{I}}$ linearly, where $\Gamma_{g_{I}}$ is defined in Corollary 1. Thus, $\beta$ is selected not to be too small to improve the invertibility of $\Gamma_{g_{I}}$. At the same time, $\beta$ should not be chosen too large, considering the controller effort, which increases with $\beta$ as seen in Eq. (22).
  3. 3.
    If the real parts of the poles of a linear system are selected as negative, then it is well known that the system is exponentially stable. In the context of ODE (30), this means that the tracking error, as well as the tracking error time derivatives, exponentially tend to zero. Poles can be adjusted so that criteria such as settling time, percentage overshoot, etc, are within a specified range. In general, poles placed further to the left on the real axis result in faster convergence but greater overshoot. We present a numerical experiment with different pole placement schemes in Example 1.
@@ 6 Numerical Experiments
$\pi(t,[x,\xi])=$ $\frac{-1}{s_{\beta}^{\prime}(\xi(t))\sqrt{16(t-0.5)^{2}-2x_{1}(t)-1}}\times\bigg(3s_{\beta}^{\prime}(\xi(t))$
$\times\sqrt{16(t-0.5)^{2}-2x_{1}(t)-1}-K\bigg[x_{1}(t),x_{2}(t),16$
$-\frac{(16t-x_{2}(t)-8)^{2}}{16(t-0.5)^{2}-2x_{1}(t)-1}-s_{\beta}(\xi(t))$
$\times\sqrt{16(t-0.5)^{2}-2x_{1}(t)-1}\bigg]^{\top}\bigg),$ (53)
@@
$u(t,[x,\xi])=$ $x_{2}+16-\frac{(16t-x_{2}-8)^{2}}{16(t-0.5)^{2}-2x_{1}-1}$ (54)
$-s_{\beta}(\xi)\sqrt{16(t-0.5)^{2}-2x_{1}-1},$
@@
To illustrate the influence of pole placement on controller performance, we performed numerical experiments using Algorithm 1 and Algorithm 2 to synthesise closed-loop controllers for different pole locations. The resulting simulations of system trajectories are shown in Fig. 3a. Each simulation applied the derived controller to the system while monitoring the switching sets from Eq. (5.3), which simplifies to Eq. (55) in this example. Whenever a change in the $\varepsilon$-NRD was detected, the controller was re-synthesised about the current state trajectory using the same procedure. The parameters were fixed as $\beta=100$ and $\varepsilon=[0.01,\dots,0.01]^{\top}$, with poles assigned either to 2.9-2.9 or 8-8.
@@
$\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\\ \dot{x}_{3}\end{bmatrix}$ $=\begin{bmatrix}10(x_{1}-x_{2})\\ 28x_{1}-x_{2}-x_{1}x_{3}\\ x_{1}x_{2}-\frac{8}{3}x_{3}\end{bmatrix}+\begin{bmatrix}0\\ 1\\ 0\end{bmatrix}u,\text{ }x(0)=\begin{bmatrix}0.1\\ 1\\ 16\end{bmatrix}$
$y(t)$ $=x_{2}(t)\text{ with }y_{r}(t)=0.6(\sin(2t)-\cos(2t))$
$\phi(t,x)$ $=[\phi_{1}(t,x),\phi_{2}(t,x)]^{\top}=[-x_{2}-1,x_{2}-1.5]^{\top}\leq 0.$
@@
Numerical experiments were conducted using an integral parameter of $\beta=100$, $\varepsilon$-NRD threshold parameter of $\varepsilon=[0.01,\dots,0.01]$ and positioning the poles at 0.3-0.3. By initialising the virtual input to have zero initial value according to Eq. (24), the initial condition of integral states was found to be $\xi(0)=[\xi_{1}(0),\xi_{2}(0)]^{\top}=[0.002,-0.001]^{\top}$. We execute Algorithm 1 and Algorithm 2 to synthesise a controller and simulate the closed-loop system. During simulation, we monitor the switching condition in Eq. (43). When triggered, we re-execute two algorithms to synthesise a new controller, switch to the updated system, and continue the simulation, again tracking the switching condition.